A health psychologist predicted that premature babies would gain more weight when they were massaged compared to being rocked. To test this prediction and with the consent of the parents, 8 premature babies in Regency Hospital participated in the experiment. All the babies were weighed before the experiment began. The babies were rocked for 20 minutes twice a day for two days and then weighed. Then the same babies were given a body massage for 20 minutes twice a day for two days and at then weighed. The change in their weight (measured in ounces) was the dependent variable.   The data are shown below.   

Baby Rocked Massaged
1 27.6 29.7
2 25.4 28.5
3 30.5 39.8
4 33.8 38.8
5 24.5 25.0
6 16.4 22.6
7 19.8 24.7
8 23.8 26.9



What is the one tail critical value at p = .05?

Select one:

a. 2.145

b. 2.120

c. 1.895

d. 2.365
Answer #1

here degree of freedom =n-1=8-1 =7

for 7 degree of freedom and 0.05 level with 1 tailed test ; critical value =1.895

option C is correct